### v1

@@ Abstract
The main purpose of this work is to present a SIMD-vectorized implementation of the symplectic 16th-order 8-stage implicit Runge-Kutta integrator based on collocation with Gauss-Legendre nodes (IRKGL16-SIMD), and to show that it can outperform state-of-the-art symplectic explicit integrators for high-precision numerical integrations (in double-precision floating-point arithmetic) of non-stiff Hamiltonian ODE systems. Our IRKGL16-SIMD integrator leverages Single Instruction Multiple Data (SIMD) based parallelism (in a way that is transparent to the user) to significantly enhance the performance of the sequential IRKGL16 implementation. We present numerical experiments comparing IRKGL16-SIMD with state-of-the-art high-order explicit symplectic methods for the numerical integration of several Hamiltonian systems in double-precision floating-point arithmetic.
@@ 1 Introduction
The family of $s$-stage implicit Runge-Kutta schemes based on collocation with Gauss-Legendre nodes (IRKGL) are known to be symmetric, symplectic and super-convergent of order $2s$, and hence are very well suited for high precision long-term numerical integration of Hamiltonian systems [18, 11]. The computational cost per step of an implicit RK method is typically higher than that of an explicit RK method, as a system of algebraic equations must be solved at each step of implicit integrators. For non-stiff problems, fixed-point iterations are commonly employed to approximately solve such systems of algebraic equations at each step. A fixed-point implementation (based on the ideas presented in [1] and written in Julia language) of the $16$th order $8$-stage IRKGL scheme (IRKGL16) is included in the DifferentialEquations.jl suite [2, 16]. Our code, IRKGL16, includes a standard time-step adaptivity mechanism. However, for long-term integration of Hamiltonian systems, the use of a constant time step is recommended [18, 11, 13]. Therefore, in this work, we will focus on its application in constant step-size mode.
@@
While there are no explicit symplectic Runge-Kutta schemes, optimized high-order explicit symplectic integrators are available for Hamiltonian systems with suitable structure. These integrators, that can be derived with the techniques of splitting and composition, are applicable when the Hamiltonian function can be decomposed into the sum of two, three, or more exactly solvable Hamiltonians [5, 11, 13].
It is widely acknowledged that high-order explicit symplectic methods are more efficient than IRKGL schemes for computations in double-precision floating-point arithmetic [11]. In fact, to achieve similar accuracy over a given integration interval, our sequential implementation of IRKGL16 usually requires more CPU time than the most effective high order explicit symplectic integrators in double-precision computations.
However, it is well established [11] that the efficiency of implicit Runge-Kutta methods can be significantly enhanced through various forms of parallelism. Single Instruction Multiple Data (SIMD), often referred to as SIMD-vectorization, is a type of parallelism that can be leveraged in most modern CPU cores, enabling vector computations at a cost comparable to scalar operations. The structure of the IRKGL16 integrator is particularly well-suited for exploiting SIMD parallelism with vectors containing eight (or alternatively, four or two) floating-point numbers.
The primary objective of this work is to present a SIMD-vectorized implementation of IRKGL16 in constant step-size mode (referred to as IRKGL16-SIMD) and to demonstrate its superiority over state-of-the-art explicit symplectic integrators for high-precision numerical integrations in double-precision floating-point arithmetic for several relevant Hamiltonian problems with a separable structure.
Our implementation makes use of the Julia package SIMD [19] to explicitly handle computations on vectors containing a fixed number of floating-point numbers. The algorithm IRKGL16-SIMD automatically performs vectorization, ensuring that the process remains transparent to the user.
@@
The integrator IRKGL16-SIMD is designed to numerically integrate non-stiff systems of (non-necessarily Hamiltonian) ODEs whose right-hand side is implemented as a generic Julia function using arithmetic operations and elementary functions supported by the SIMD package.
@@
The paper is organized as follows: Section 1 provides an introduction; Section 2 describes a vectorized fixed-point implementation of symplectic implicit Runge-Kutta (IRK) schemes; Section 3 presents implementation details of the IRKGL16-SIMD method; Section 4 reviews state-of-the-art symplectic explicit integrators; Section 5 reports numerical experiments; and Section 6 presents the conclusions.
@@ 2.1 Implicit RK methods
We consider initial value problems for systems of ordinary differential equations (ODEs) of the form:
@@
[…]
The RK approximations $y_{n}\approx y(t_{n})$ to the solution $y(t)$ at times $t_{n}=t_{0}+nh$ for $n=1,2,3,\dots$ are given by:
@@
Here, each $Y_{n,i}$ is an approximation of the state vector $y(t)$ at the intermediate time $t=t_{n-1}+hc_{i}$ computed within each time-step. The coefficients $a_{ij}$, $b_{i}$, and $c_{i}$ from the Butcher tableau determine the specific RK method and its properties.
If the matrix $A=(a_{ij})$ is lower triangular the Runge-Kutta method is explicit. Otherwise, it is an implicit Runge-Kutta (IRK) scheme, and the stage vectors $Y_{n,i}$ at each step are defined implicitly by the equation (3).
@@ 2.2 IRK methods of collocation type
In a collocation method based on the nodes $c_{1},\ldots,c_{s}\in[0,1]$, the approximations $y_{n}\approx y(t_{n})$ to the solution $y(t)$ at times $t_{n}=t_{0}+nh$ are obtained as follows: for each $n=1,2,3,\dots$, consider the polynomial function $P_{n}:\mathbb{R}\to\mathbb{R}^{D}$ of degree $s$ uniquely determined by the following conditions,
@@
This is equivalent [11] to the application of the IRK method (2)–(3) with coefficients $b_{i},a_{ij}$ uniquely determined by the following conditions:
@@
In the specific case where $c_{1},\ldots,c_{s}\in(0,1)$ are the shifted Gauss-Legendre nodes–defined by $c_{i}=(1+x_{i})/2$, where $x_{1},\ldots,x_{s}$ are the zeroes of the Legendre polynomial of degree ss–the resulting IRK scheme of collocation type, which we refer to as the $s$-stage IRKGL scheme, achieves an order of accuracy $2s$ [11].
2.3 Symplectic IRK methods with floating-point coefficients
We consider an autonomous Hamiltonian system of the form
$\frac{dy}{dt}=J^{-1}\nabla H(y),\quad y\in\mathbb{R}^{2d},$ (5)
$J=\begin{pmatrix}0&I_{d}\\ -I_{d}&0\end{pmatrix}$.
The exact flow of this system, denoted by $\varphi_{H}^{t}$, is symplectic, meaning that it preserves the symplectic form: for all $t$, the Jacobian matrix $D\varphi_{H}^{t}(y)$ satisfies
$\left(D\varphi_{H}^{t}(y)\right)^{T}JD\varphi_{H}^{t}(y)=J.$
$b_{i}a_{ij}+b_{j}a_{ji}=b_{i}b_{j},\quad 1\leqslant i,j\leqslant s.$ (6)
As shown in [12], if the coefficients $b_{i},a_{ij}$ are replaced by machine numbers $\tilde{b}_{i},\tilde{a}_{ij}$, then the resulting implicit RK scheme typically fails to satisfy the simplectic condition (6). To ensure that the symplectic property is exactly preserved with double-precision floating-point coefficients, we reformulate the IRKGL schemes as follows [1]:
$y_{n}$ $=y_{n-1}+\sum^{s}_{i=1}L_{n,i},$ (8)
where $\mu_{ij}=a_{ij}/b_{j},\quad 1\leqslant i,j\leqslant s$. If the coefficients $b_{i}$, $c_{i}$, and $\mu_{ij}$ are replaced by the double-precision floating-point coefficients $\tilde{b}_{i}$, $\tilde{c}_{i}$, and $\tilde{\mu}_{ij}$, then the resulting scheme is still symplectic provided that $\tilde{\mu}_{ij}$ is set for $1\leqslant i<j\leqslant s$ as $1-\tilde{\mu}_{ji}$.
2.4 Vectorized fixed-point iteration
The primary challenge in implementing implicit Runge-Kutta (IRK) methods is the efficient solution of the nonlinear system for the internal stages. For non-stiff problems, fixed-point iteration is generally recommended. In collocation-type IRK methods, it is standard practice to initialize the fixed-point iteration with an estimate of the internal stages extrapolated from the previous time-step, that is,
@@ 2.4 Vectorized fixed-point iteration
We introduce notation to facilitate the description of the proposed vectorized implementation of the fixed-point iteration.
Let $b$ and $c$ denote the $s$-vectors $(b_{1},\ldots,b_{s})$ and $(c_{1},\ldots,c_{s})$, respectively. Similarly, for each $i=1,\ldots,s$, let $\mu_{i}$ and $\nu_{i}$ represent the $s$-vectors $(\mu_{1i},\ldots,\mu_{si})$ and $(\nu_{1i},\ldots,\nu_{si})$, respectively.
For an arbitrary $s$-vector $v=(v_{1},\ldots,v_{s})$ and a real number $\lambda\in\mathbb{R}$, we denote by $\lambda+v$ the $s$-vector $(\lambda+v_{1},\ldots,\lambda+v_{s})$. Given two arbitrary $s$-vectors $v=(v_{1},\ldots,v_{s})$ and $w=(w_{1},\ldots,w_{s})$, $v\circ w$ represents the componentwise product of two $s$-vectors defined by $(v_{1}w_{1},\ldots,v_{s}w_{s})$. We also use the notation $\mathrm{sum}(v)=v_{1}+\cdots+v_{s}$ and $\|v\|_{\infty}=\max(|v_{1}|,\ldots,|v_{s}|)$.
For each $j\in\{1,\ldots,D\}$, we denote by $Y_{n,i}^{j}$, $L_{n,i}^{j}$, and $f^{j}(t_{n-1}+hc_{i},Y_{n,i})$ the $j$th component of $Y_{n,i}\in\mathbb{R}^{D}$, $L_{n,i}\in\mathbb{R}^{D}$ and $f(t_{n-1}+hc_{i},Y_{n,i})\in\mathbb{R}^{D}$ respectively. For each $j\in\{1,\ldots,D\}$, we consider the $s$-vectors $Y_{n}^{j}=(Y_{n,1}^{j},\ldots,Y_{n,s}^{j})$ and $L_{n}^{j}=(L_{n,1}^{j},\ldots,L_{n,s}^{j})$.
Let $Y_{n}$ represent the one-dimensional array $(Y_{n}^{1},\ldots,Y_{n}^{D})$ of $s$-vectors $Y_{n}^{j}$. We define the function $F$ that given $t_{n-1}+hc$ and $Y_{n}$, returns the one-dimensional array $(F_{n}^{1},\ldots,F_{n}^{D})$ of $s$-vectors
$F_{n}^{j}=(f^{j}(t_{n-1}+hc_{1},Y_{n,1}),\ldots,f^{j}(t_{n-1}+hc_{s},Y_{n,s})).$
@@
$\begin{split}F_{n}&=F(t_{n-1}+hc,Y_{n}),\\ L_{n}^{j}&=h\,(b\circ F_{n}^{j}),\quad j=1,\ldots,D,\\ Y_{n}^{j}&=y_{n-1}^{j}+\sum_{i=1}^{s}\mu_{i}L_{n,i}^{j},\quad j=1,\ldots,D,\end{split}$ (11)
@@
$y_{n}^{j}=y_{n-1}^{j}+\mathrm{sum}(L_{n}^{j}),\quad j=1,\ldots,D.$ (12)

### v2

@@ Abstract
Implicit Runge-Kutta schemes based on collocation with Gauss-Legendre nodes (IRKGL) possess attractive theoretical properties for long-term integration of Hamiltonian systems—they are symmetric, symplectic, and super-convergent—yet they have traditionally been considered less practical than explicit symplectic integrators. In this work, we challenge this conventional wisdom by introducing a stage-wise SIMD-vectorization approach that enables IRKGL schemes to outperform state-of-the-art explicit symplectic integrators for high-precision computations in double-precision floating-point arithmetic. Our approach reformulates the fixed-point iteration in terms of $s$-vectors (where $s$ is the number of stages) to explicitly exploit Single Instruction Multiple Data (SIMD) capabilities of modern processors, significantly reducing computational overhead and enabling parallel evaluation of the right-hand side function across all stages. Additionally, we present a reformulation of IRKGL schemes for second-order ODEs that ensures exact symplecticity at the level of double-precision floating-point arithmetic, extending our previous approach for first-order systems to this setting. We demonstrate these ideas through IRKGL16, a Julia implementation of the 16th-order 8-stage IRKGL scheme that performs vectorization seamlessly and transparently to the user. Numerical experiments on several Hamiltonian problems with separable structure confirm the effectiveness of our approach.
@@ 1 Introduction
The family of $s$-stage implicit Runge-Kutta schemes based on collocation with Gauss-Legendre nodes (IRKGL) are known to be symmetric, symplectic and super-convergent of order $2s$, making them theoretically well-suited for high-precision long-term numerical integration of Hamiltonian systems [21, 13]. However, despite these attractive theoretical properties, IRKGL schemes have traditionally been considered less practical than explicit symplectic integrators for Hamiltonian problems that admit a separable structure. The main reasons for this preference include: (i) the need for careful implementation to avoid linear drift of energy error [14, 1], (ii) the higher computational overhead inherent to implicit methods, and (iii) the typically larger number of right-hand side evaluations required to achieve comparable accuracy.
@@
For Hamiltonian systems with suitable structure—where the Hamiltonian function can be decomposed into the sum of two or more exactly solvable components—optimized high-order explicit symplectic integrators derived through splitting and composition techniques [8, 6, 13, 16] have thus become the methods of choice for high-precision integration. These methods avoid the complications associated with implicit schemes while maintaining the crucial symplectic property.
In this work, we challenge this conventional wisdom by showing that IRKGL schemes can outperform state-of-the-art explicit symplectic integrators for high-precision computations in double-precision floating-point arithmetic. Our main contribution is a novel approach to formulating and implementing the fixed-point iterations of IRKGL schemes: we reformulate the iterations in terms of $s$-dimensional vectors (where $s$ is the number of stages) and design the algorithm to explicitly exploit Single Instruction Multiple Data (SIMD) capabilities of modern processors.
This stage-wise SIMD-vectorization approach represents a departure from conventional strategies for reducing overhead in implicit integrators. Traditionally, SIMD-vectorization is achieved by structuring inner loops in a way that allows the compiler to automatically vectorize them. In contrast, our approach explicitly builds SIMD operations into the algorithm design itself. This has two crucial advantages: it significantly reduces the computational overhead of the implicit method, and it enables parallel evaluation of the right-hand side function across all $s$ stages simultaneously (provided the function satisfies certain requirements for vectorization).
To ensure exact symplecticity at the level of double-precision floating-point arithmetic, we address the linear drift in energy error that can affect symplectic IRK schemes [14]. For first-order ODEs, we adopt the approach presented in [1]. Additionally, we generalize this approach to the implicit Runge-Kutta-Nyström (IRKN) formulation for second-order ODEs, which is the natural framework for many Hamiltonian systems arising from classical mechanics.
We demonstrate these ideas through a concrete implementation, IRKGL16, which applies our stage-wise SIMD-vectorization strategy to the $16$th-order 88-stage IRKGL scheme (i.e., $s=8$, order $2s=16$). The implementation, written in Julia using the SIMD package [22], performs vectorization automatically and seamlessly, remaining transparent to the user. While IRKGL16 includes both constant step-size and adaptive step-size modes, in this work we focus on the constant step-size mode, as recommended for long-term symplectic integration of Hamiltonian systems [21, 13, 16]. We have also implemented IRKGL methods with $s=2$, $s=4$, and $s=6$ stages, which are available in the IRKGL16 code, though our numerical experiments (see Subsection 5.2) demonstrate that the $8$-stage method achieves the best performance for high-precision computations.
@@
The code is designed for non-stiff systems of ODEs (not necessarily Hamiltonian). While it can handle arbitrary precision arithmetic, the vectorized implementation is available for single- and double-precision floating-point arithmetic. Maximum efficiency through full vectorization is achieved when the right-hand side function is implemented as a generic Julia function using arithmetic operations and elementary functions supported by the SIMD package.
Through numerical experiments on several representative Hamiltonian problems with separable structure, we show that our stage-wise SIMD-vectorization approach enables IRKGL16 to outperform state-of-the-art explicit symplectic integrators for high-precision integration tasks in double-precision arithmetic.
@@
The paper is organized as follows: Section 2 describes the reformulation of IRKGL schemes for both first-order and second-order ODEs to ensure exact symplecticity and our stage-wise SIMD-vectorization strategy; Section 3 presents implementation details of IRKGL16; Section 4 reviews state-of-the-art high-order explicit symplectic integrators; Section 5 reports numerical experiments; and Section 6 presents the conclusions.
@@ 2.1 Implicit RK methods
In this work, we focus on the numerical integration of autonomous Hamiltonian systems of the form
$\frac{dy}{dt}=J^{-1}\nabla H(y),\quad y\in\mathbb{R}^{2d},$ (1)
$J=\begin{pmatrix}0&I_{d}\\ -I_{d}&0\end{pmatrix}$.
[…]
Although our focus is on Hamiltonian systems, RK methods are applicable to more general ODE problems of the more general form:
@@
The RK approximations $y_{n}\approx y(t_{n})$ to the solution $y(t)$ of (2) at times $t=t_{n}=t_{0}+nh$ for $n=1,2,3,\dots$ are given by:
@@
Here, each $Y_{n,i}$ is an approximation of the state vector $y(t)$ at the intermediate time $t=t_{n-1}+hc_{i}$ computed within each time-step. The vectors $b=(b_{1},\ldots,b_{s})$, $c=(c_{1},\ldots,c_{s})$, and the matrix $A=(a_{ij})$ from the Butcher tableau determine the specific RK method and its properties.
If the matrix $A$ is lower triangular the Runge-Kutta method is explicit. Otherwise, it is an implicit Runge-Kutta (IRK) scheme, and the stage vectors $Y_{n,i}$ at each step are defined implicitly by the equation (4).
The primary challenge in implementing implicit Runge-Kutta (IRK) methods is the efficient solution of the nonlinear system for the internal stages. For non-stiff problems, fixed-point iteration is generally recommended [21, 13].
@@ 2.2 IRK methods of collocation type
Among IRK methods, those based on collocation are particularly attractive. In a collocation method based on the pairwise distinct nodes $c_{1},\ldots,c_{s}\in[0,1]$, the approximations $y_{n}\approx y(t_{n})$ to the solution $y(t)$ at times $t_{n}=t_{0}+nh$ are obtained as follows: for each $n=1,2,3,\dots$, consider the polynomial function $P_{n}:\mathbb{R}\to\mathbb{R}^{D}$ of degree $s$ uniquely determined by the following conditions,
@@
This collocation procedure is equivalent [13] to the application of the IRK method (3)–(4) with coefficients $b_{i},a_{ij}$ uniquely determined by the following conditions:
@@
In the specific case where $c_{1},\ldots,c_{s}\in(0,1)$ are the shifted Gauss-Legendre nodes (defined by $c_{i}=(1+x_{i})/2$, where $x_{1},\ldots,x_{s}$ are the zeros of the Legendre polynomial of degree $s$) the resulting IRK scheme of collocation type, which we refer to as the $s$-stage IRKGL scheme, achieves an order of accuracy $2s$ [13].
In collocation-type IRK methods, it is standard practice to initialize the fixed point iteration with an estimate of the internal stages extrapolated from the previous time-step, that is,
@@
2.3 Application of IRK methods to second order differential equations
Autonomous Hamiltonian systems with a Hamiltonian function of the form
$H(q,p)=\frac{1}{2}\,p^{T}M^{-1}p+U(q)$ (8)
can be expressed as system of second order ODEs of the form
$\frac{d^{2}}{dt^{2}}q=M^{-1}\nabla U(q).$ (9)
Any RK scheme can be applied to systems of second-order differential equations $\frac{d^{2}}{dt^{2}}q=g(t,q)$ by rewriting it as a system of first-order differential equations,
$\frac{dq}{dt}=v,\quad\frac{dv}{dt}=g(q,t),$ (10)
Obviously, the system (10) is a specific instance of the general form (2), with $D=2d$, $y=(y^{1},\ldots,y^{2d})$, $q=(y^{1},\ldots,y^{d})$, $v=(y^{d+1},\ldots,y^{2d})$, and $f(t,y)=(v,g(t,q))$.
It is well known that the implicit equations of the IRK scheme can be simplified in the case of second order ODE systems of the form (10) by rewriting it as an implicit Runge-Kutta-Nystøm (IRKN) method. More precisely, the approximations $y_{n}=(q_{n},v_{n})\approx(q(t_{n}),v(t_{n}))$ of the solution $y(t)=(q(t),v(t))$ of the (10) supplemented with the initial conditions
$q(t_{0})=q_{0},\quad v(t_{0})=v_{0}$ (11)
can be computed as follows: for $n=1,2,\ldots$
$\begin{split}q_{n}&=q_{n-1}+h\,v_{n-1}+h^{2}\sum_{i=1}^{s}\beta_{i}\,g(t_{n-1}+c_{i}h,Q_{n,i}),\\ v_{n}&=v_{n-1}+h\sum_{i=1}^{s}b_{i}g(t_{n-1}+c_{i}h,Q_{n,i}),\end{split}$ (12)
where the vectors $Q_{n,i}$, $i=1,\ldots,s$, are implicitly defined by
$Q_{n,i}=q_{n-1}+h\,c_{i}\,v_{n-1}+h^{2}\,\sum_{j=1}^{s}\alpha_{ij}\,g(t_{n-1}+c_{j}h,Q_{n,j}),\quad i=1,\ldots,s.$ (13)
Here, $\beta_{i}=\sum_{k=1}^{s}b_{k}a_{ki}$, $c_{i}=\sum_{k=1}^{s}a_{ik}$, and $\alpha_{ij}=\sum_{k=1}^{s}a_{ik}a_{kj}$. This formulation gives rise to a more efficient fixed-point iteration, which, compared to the standard IRK formulation, approximately halves the required number of iterations for solving the implicit equations [21, 13].
2.4 Symplectic IRK methods with floating-point coefficients
Symplectic IRK methods
The $t$-flow of the Hamiltonian system (1), denoted by $\varphi_{t}$, is symplectic, meaning that it preserves the symplectic form: for all $t$, the Jacobian matrix $D\varphi_{t}(y)$ satisfies
$\left(D\varphi_{t}(y)\right)^{T}JD\varphi_{t}(y)=J.$
$b_{i}a_{ij}+b_{j}a_{ji}=b_{i}b_{j},\quad 1\leqslant i,j\leqslant s.$ (14)
Symplectic IRKN methods
An IRK method rewritten as an IRKN (12)–(13) is symplectic if its coefficients satisfy the conditions [25, 21]
$\begin{split}\beta_{i}&=b_{i}\,(1-c_{i}),\quad 1\leqslant i\leqslant s,\\ b_{i}\,(\beta_{j}-\alpha_{ij})&=b_{j}\,(\beta_{i}-\alpha_{ji}),\quad 1\leqslant i,j\leqslant s.\end{split}$ (15)
Reformulation for floating-point arithmetic
As shown in [14], if the coefficients $b_{i},a_{ij}$ of the IRK schemes (or $b_{i},\beta_{i},c_{i},\alpha_{ij}$ of the IRKN formulation) are replaced by their machine number representatives, then the resulting implicit RK scheme typically fails to satisfy the symplecticity condition (14). To ensure that the symplectic property is exactly preserved with double-precision floating-point coefficients, for first-order systems, we adopt the reformulation from [1]:
$y_{n}$ $=y_{n-1}+\sum^{s}_{i=1}L_{n,i},$ (17)
where $b_{i}$ and $c_{i}$ are the machine-number representatives of the original coefficients of the IRK scheme, and
$\mu_{ij}$ $=\mathrm{fl}(a_{ij}/b_{j}),\quad 1\leqslant j\leqslant i\leqslant s,$
$\mu_{ij}$ $=1-\mu_{ji},\quad 1\leqslant i<j\leqslant s.$
Here, $\mathrm{fl}(x)$ represents the machine-number representative of the real number $x$.
For second-order systems, we generalize this approach as follows: We reformulate the IRKN form (12)–(13) of IRKGL schemes in terms of the coefficients $b_{i}$, $c_{i}$, and $\eta_{ij}=\alpha_{ij}/b_{j}$, so that the second condition in (15) can be replaced by
$\eta_{ij}+c_{j}=\eta_{ji}+c_{i}.$ (18)
Now, we define one step of the method as follows:
$G_{n,i}$ $=g(t_{n-1}+hc_{i},Q_{n,i}),\quad R_{n,i}=h\,b_{i}\,G_{n,i},\quad i=1,\ldots,s,$ (19)
$Q_{n,i}$ $=q_{n-1}+h\,c_{i}\,v_{n-1}+h\sum^{s}_{j=1}\eta_{ij}\,R_{n,j},\quad i=1,\ldots,s,$ (20)
$v_{n}$ $=v_{n-1}+\sum^{s}_{i=1}R_{n,i},\quad q_{n}=q_{n-1}+h\,v_{n}-h\sum^{s}_{i=1}c_{i}R_{n,i}$ (21)
where the coefficients $b_{i},c_{i}$ are machine-number representatives of the original coefficients of the $s$-stage IRKGL scheme, and
$\eta_{ij}$ $=\mathrm{fl}(\alpha_{ij}/b_{j}),\quad 1\leqslant j\leqslant i\leqslant s,$
$\eta_{ij}$ $=c_{i}-c_{j}-\eta_{ji},\quad 1\leqslant i<j\leqslant s.$
This guarantees that the symplecticity conditions (18) holds. We have verified numerically that the coefficients $\eta_{ij}$ computed by this procedure are exactly representable as machine numbers in double-precision arithmetic for $s\leqslant 16$.
2.5 Stage-wise vectorization of fixed-point iteration
In order to enable SIMD implementation, we reformulate the method using $s$-vectors that group values across all stages.
@@ 2.5 Stage-wise vectorization of fixed-point iteration
We introduce notation to facilitate the description of the proposed stage-wise vectorized implementation of the fixed-point iteration.
Let $\boldsymbol{b}$ and $\boldsymbol{c}$ denote the $s$-vectors $(b_{1},\ldots,b_{s})$ and $(c_{1},\ldots,c_{s})$, respectively. Similarly, for each $i=1,\ldots,s$, let $\boldsymbol{\mu_{i}}$, $\boldsymbol{\nu_{i}}$, and $\boldsymbol{\eta_{i}}$ represent the $s$-vectors $(\mu_{1i},\ldots,\mu_{si})$, $(\nu_{1i},\ldots,\nu_{si})$, and $(\eta_{1i},\ldots,\eta_{si})$, respectively.
For an arbitrary $s$-vector $\boldsymbol{v}=(v_{1},\ldots,v_{s})$ and a real number $\lambda\in\mathbb{R}$, we denote by $\lambda+\boldsymbol{v}$ the $s$-vector $(\lambda+v_{1},\ldots,\lambda+v_{s})$. Given two arbitrary $s$-vectors $\boldsymbol{v}=(v_{1},\ldots,v_{s})$ and $\boldsymbol{w}=(w_{1},\ldots,w_{s})$, $\boldsymbol{v}*\boldsymbol{w}$ represents the componentwise product of two $s$-vectors defined by $(v_{1}w_{1},\ldots,v_{s}w_{s})$. We also use the notation $\mathrm{sum}(\boldsymbol{v})=v_{1}+\cdots+v_{s}$, and $\|\boldsymbol{v}\|_{\infty}=\max(|v_{1}|,\ldots,|v_{s}|)$.
To distinguish $s$-vectors from collections of $D$ state variables, we denote the laters as
$y=\left[\begin{matrix}y^{1}\\ \vdots\\ y^{D}\end{matrix}\right]$
(or alternatively as $y=[y^{1},\ldots,y^{D}]$) and refer to them as arrays of state variables.
For each $j\in\{1,\ldots,D\}$, we denote by $Y_{n,i}^{j}$, $L_{n,i}^{j}$, and $f^{j}(t_{n-1}+hc_{i},Y_{n,i})$ the $j$th component of the arrays $Y_{n,i}\in\mathbb{R}^{D}$, $L_{n,i}\in\mathbb{R}^{D}$ and $f(t_{n-1}+hc_{i},Y_{n,i})\in\mathbb{R}^{D}$ respectively. For each $j\in\{1,\ldots,D\}$, we consider the $s$-vectors $\boldsymbol{Y_{n}^{j}}=(Y_{n,1}^{j},\ldots,Y_{n,s}^{j})$ and $\boldsymbol{L_{n}^{j}}=(L_{n,1}^{j},\ldots,L_{n,s}^{j})$.
Let $\boldsymbol{Y_{n}}$ and $\boldsymbol{F_{n}}$ represent the arrays of $s$-vectors $\boldsymbol{Y_{n}^{j}}$ and $\boldsymbol{F_{n}^{j}}$
$\boldsymbol{Y_{n}}=\left[\begin{matrix}\boldsymbol{Y_{n}^{1}}\\ \vdots\\ \boldsymbol{Y_{n}^{D}}\end{matrix}\right],\quad\boldsymbol{F_{n}}=\left[\begin{matrix}\boldsymbol{F_{n}^{1}}\\ \vdots\\ \boldsymbol{F_{n}^{D}}\end{matrix}\right].$
We define the vectorized function $\boldsymbol{f}$ that given $t_{n-1}+h\boldsymbol{c}$ and $\boldsymbol{Y_{n}}$, returns the array $\boldsymbol{F_{n}}$ of $s$-vectors
$\boldsymbol{F_{n}^{j}}=(f^{j}(t_{n-1}+hc_{1},Y_{n,1}),\ldots,f^{j}(t_{n-1}+hc_{s},Y_{n,s})),\quad j=1,\ldots,D.$
That function $\boldsymbol{f}$ represents the parallel evaluation of $f(t_{n-1}+hc_{1},Y_{n,1}),\ldots,f(t_{n-1}+hc_{s},Y_{n,s})$. For instance, if $D=2$, $s=2$, and $f(t,[y^{1},y^{2}])=[2\,y^{1}*y^{2},t*(y^{1}+y^{2})]$, then $\boldsymbol{f}(t+h\boldsymbol{c},\boldsymbol{Y_{n}})$ returns the array
$\left[\begin{matrix}2\,\boldsymbol{Y_{n}}^{1}*\boldsymbol{Y_{n}}^{2}\\ (t+h\boldsymbol{c})*(\boldsymbol{Y_{n}}^{1}+\boldsymbol{Y_{n}}^{2})\end{matrix}\right]=\left[\begin{matrix}(2\,Y_{n,1}^{1}*Y_{n,1}^{2},Y_{n,2}^{1}*Y_{n,2}^{2})\\ ((t+hc_{1})*(Y_{n,1}^{1}+Y_{n,1}^{2}),(t+hc_{2})*(Y_{n,2}^{1}+Y_{n,2}^{2}))\end{matrix}\right]$
Note how operations on $s$-vectors are performed component-wise, enabling parallel evaluation across all stages.
@@
$\begin{split}\boldsymbol{F_{n}}&=\boldsymbol{f}(t_{n-1}+h\,\boldsymbol{c},\boldsymbol{Y_{n}}),\\ \boldsymbol{L_{n}^{j}}&=h\,(\boldsymbol{b}*\boldsymbol{F_{n}^{j}}),\quad j=1,\ldots,D,\\ \boldsymbol{Y_{n}^{j}}&=y_{n-1}^{j}+\sum_{i=1}^{s}\boldsymbol{\mu_{i}}\,L_{n,i}^{j},\quad j=1,\ldots,D,\end{split}$ (22)
@@
$y_{n}^{j}=y_{n-1}^{j}+\mathrm{sum}(\boldsymbol{L_{n}^{j}}),\quad j=1,\ldots,D.$ (23)